A downtown bank is advertising that if you deposit $1,000 with them, and leave it there for 65 months, you can get $2,000 back at the end of this period. Assuming monthly compounding, what is the approximate monthly rate of interest paid by the bank?
Answer #1

F = 2000, P = 1000, T = 65 months

F = P *(1+i)^t

2000 = 1000 * (1+i)^65

(1+i)^65 = 2000 / 1000 = 2

1+i = 2 ^(1/65) = 1.01072

i = 1.01072 - 1 = 0.01072 ~ 1.07% per month
Answer #2

F = P *(1+i)^t

 

2000 = 1000 * (1+i)^65

 

(1+i)^65 = 2000 / 1000 = 2

 

1+i = 2 ^(1/65)

 

i = 1.01072

 

i = 1.01072 - 1 = 0.01072 or 1.07% per month


answered by: Rannie Asis